OpenStudy (anonymous):

Give a big-O estimate for the number additions used in this segment of an algorithm. t := 0 for i := 1 to n for j := 1 to n t := t + i + j

OpenStudy (anonymous):

The composition of functions has their big-O estimates multiplied. In this case: \(O(n\times n\times 1) = O(n^2)\)